OpenStudy (anonymous):

Indicate the equation of the line meeting the given conditions. Please put the equation in standard form. Containing A(1, 3) and B(0, 2)

OpenStudy (anonymous):

m= (y1-y)/(x1-x) m=(3-2)/(1-0)=1 then the equation of the line is: y-2=1*(x-0) y=x+2